OpenStudy (anonymous):

How many solutions does the equation –3a + 3a + 6 = 7 have? (5 points) Infinitely many Two One None

OpenStudy (anonymous):

What you think it is?

OpenStudy (anonymous):

I think it's One solution. What do you think?

OpenStudy (anonymous):

i thought it was none but idk

OpenStudy (diamondboy):

solve for a and see how many solutions it has

OpenStudy (anonymous):

this is why i think one.. first you..Rearrange the equation by subtracting what is to the right of the equal sign from both sides of the equation : 3*a+3*a+6-(7)=0 Solve:6a-1 = 0 Add 1 to both sides of the equation :6a = 1 Divide both sides of the equation by 6: a = 1/6

OpenStudy (diamondboy):

u forgot a d -ve sign @amandaapandaa_7

OpenStudy (anonymous):

oh so it is one

OpenStudy (diamondboy):

yep

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

Sorry forgot it. LOl.